# Irreversibility in scalar active turbulence: The role of topological defects

### v3

@@ 4 Discussion
Using the formalism of stochastic thermodynamics [48, 31, 50], we demonstrate that the dominant contribution to IEPR is generically located in regions of high vorticity. This contribution is due to the coupling between the momentum-conserving fluid (governed by the Stokes equation) and the density of swimmers through active stress. Moreover, using linear irreversible thermodynamics, we have shown that the IEPR represents only a partial estimation of the total dissipation rate. To analyze in detail the spatial profile of IEPR, we build a mapping between the shear stress, which stirs the fluid, and an effective nematic field, by analogy with liquid crystals [12]. Our mapping leads to identifying some $\pm\frac{1}{2}$ defects, with symmetries analogous to that of nematics, defined as locations where the gradient of the swimmer density vanishes. Similarly to active nematics, topological defects stir the flow and shape the spatial structure of vortices. Here, we have shown that, remarkably, the vortices near topological defects indicate the dominant contribution to irreversibility. Therefore, our mapping reveals a direct relation between topological defects and IEPR.
We provide analytical derivations for the profile of IEPR around $\frac{1}{2}$ and $\frac{1}{2}$ defects. Our results predict that the dominant contribution to IEPR stems from configurations where nearby $\frac{1}{2}$ defects are oriented orthogonally to each other, in direct agreement with numerical simulations. Moreover, we examine the dependence of IEPR on the distance between $(+\frac{1}{2},-\frac{1}{2})$ pairs, as a first step towards quantifying IEPR during the creation/annihilation of such pairs (a detailed analysis requires resolving the defect dynamics during creation/annihilation) [77]. These results open some unprecedented perspectives for quantifying irreversibility in experimental realizations of active turbulence. Moreover, the strong localization of IEPR in small regions centered around the point-like defects indicates that the total IEPR across the system may be estimated by a discrete number of local measurements. It would be interesting to explore further whether the fluctuations of IEPR are also constrained by the defect statistics [78].
@@ Appendix A Stochastic dynamics of the stream function
where we have used the identities $\epsilon_{\alpha\alpha^{\prime}}\epsilon_{\alpha\mu^{\prime}}=\delta_{\alpha^{\prime}\mu^{\prime}}$ and $\epsilon_{\alpha\alpha^{\prime}}\epsilon_{\mu\mu^{\prime}}=\delta_{\alpha\mu}\delta_{\alpha^{\prime}\mu^{\prime}}-\delta_{\alpha\alpha^{\prime}}\delta_{\mu\mu^{\prime}}$. Therefore, combining Eqs. (36-41), we deduce the stream dynamics in Eq. (9).
@@ Appendix B Numerical methods
To integrate the hydrodynamic equations [Eqs. 1 and 9], we use a pseudo-spectral numerical scheme. The simulations are performed in a square lattice of size $L\times L$, and the space is discretized as $L=N\Delta$. The grid spacing $\Delta$ is kept constant while varying the $N$ to change the system size. The initial condition for $\phi$ is sampled from randomly in [0.1,0.1][-0.1,0.1], and $\psi$ is set to zero.
@@
Figure 7: (a) EPR is plotted as a function of activity for various spatial discretizations by fixing $L$. (b) Variation of EPR with spatial discretizations. The inset plot shows EPR per lattice cell. (c) Plot is same as (a) but by changing $L$ with fixed $\Delta$.
@@ Appendix D Derivation of the informatic entropy production rate
$\dot{\cal S}_{\psi}=-\frac{1}{T}\int\langle L\psi\rangle\,\mathrm{d}{\bf\it r},\quad\dot{\cal S}_{\phi}=-\frac{1}{T}\int\left\langle\left(\dot{\phi}+{\bf\it v}\cdot\nabla\phi\right)\frac{\delta\cal F}{\delta\phi}\right\rangle\,\mathrm{d}{\bf\it r},$ (64)
@@
$T\dot{\cal S}_{\phi}$ $=\epsilon_{\alpha\beta}\int\langle(\partial_{\beta}\psi)(\partial_{\alpha}\phi)(dg/d\phi-\kappa\nabla^{2}\phi)\rangle\,\mathrm{d}{\bf\it r}$ (66)
$=\epsilon_{\alpha\beta}\int\langle(\partial_{\beta}\psi)(\partial_{\alpha}g-\kappa(\partial_{\alpha}\phi)\nabla^{2}\phi)\rangle\,\mathrm{d}{\bf\it r}$
$=-\epsilon_{\alpha\beta}\int\langle\psi(\partial_{\beta}\partial_{\alpha}g-\kappa\partial_{\beta}((\partial_{\alpha}\phi)\nabla^{2}\phi))\rangle\,\mathrm{d}{\bf\it r},$
where we have integrated by parts. Given that $\epsilon_{\alpha\alpha}\partial_{\alpha}\partial_{\beta}h(\phi)=0$ for an arbitrary function $h$ (since $\epsilon_{\alpha\beta}=-\epsilon_{\alpha\beta}$), we deduce
$T\dot{\cal S}_{\phi}=\kappa\epsilon_{\alpha\beta}\int\langle\psi(\partial_{\alpha}\phi)\nabla^{2}(\partial_{\beta}\phi)\rangle\,\mathrm{d}{\bf\it r}.$ (67)
@@
$T\dot{\cal S}_{\psi}=-(\zeta+\kappa)\epsilon_{\alpha\beta}\int\langle\psi(\partial_{\alpha}\phi)\nabla^{2}(\partial_{\beta}\phi)\rangle\,\mathrm{d}{\bf\it r}.$ (68)

### v4

@@ 4 Discussion
Using the formalism of stochastic thermodynamics [48, 31, 50], we demonstrate that the dominant contribution to IEPR is generically located in regions of high vorticity. This contribution is due to the coupling between the momentum-conserving fluid (governed by the Stokes equation) and the density of swimmers through active stress. Moreover, using linear irreversible thermodynamics, we have shown that the IEPR represents only a partial estimation of the total dissipation rate. To analyze in detail the spatial profile of IEPR, we build a mapping between the shear stress, which stirs the fluid, and an effective nematic field, by analogy with liquid crystals [12]. Our mapping leads to identifying some $\pm\frac{1}{2}$ defects, with symmetries analogous to that of nematics, defined as locations where the gradient of the swimmer density vanishes. Similarly to active nematics, topological defects stir the flow and shape the spatial structure of vortices. Here, we have shown that, remarkably, the vortices near pairs of topological defects indicate the dominant contribution to irreversibility. Therefore, our mapping reveals a direct relation between topological defects and IEPR.
We provide analytical derivations for the profile of IEPR around $\frac{1}{2}$ and $\frac{1}{2}$ defects. Our results predict that the dominant contribution to IEPR stems from configurations where nearby $\frac{1}{2}$ defects are oriented orthogonally to each other, in direct agreement with numerical simulations. Moreover, we examine the dependence of IEPR on the distance between $(+\frac{1}{2},-\frac{1}{2})$ pairs, as a first step towards quantifying IEPR during the creation/annihilation of such pairs (a detailed analysis requires resolving the TDs dynamics during creation/annihilation) [77]. We show that the main features of the local IEPR can be predicted analytically by considering point-like idealized TDs, both isolated and paired. Taking into account a finite core size for TDs, we can even predict the finer details of the local IEPR. These results open some unprecedented perspectives for quantifying irreversibility in experimental realizations of active turbulence. Moreover, the strong localization of IEPR in small regions centered around the point-like TDs indicates that the total IEPR across the system may be estimated by a discrete number of local measurements. It would be interesting to explore further whether the fluctuations of IEPR are also constrained by the TDs statistics [78].
@@ Appendix A Stochastic dynamics of the stream function
where we used the identities $\epsilon_{\alpha\alpha^{\prime}}\epsilon_{\alpha\mu^{\prime}}=\delta_{\alpha^{\prime}\mu^{\prime}}$ and $\epsilon_{\alpha\alpha^{\prime}}\epsilon_{\mu\mu^{\prime}}=\delta_{\alpha\mu}\delta_{\alpha^{\prime}\mu^{\prime}}-\delta_{\alpha\alpha^{\prime}}\delta_{\mu\mu^{\prime}}$. Therefore, combining Eqs. (38-43), we deduce the stream function’s dynamics in Eq. (10).
@@ Appendix B Numerical methods
To integrate the hydrodynamic equations [Eqs. 1 and 10], we use a pseudo-spectral numerical scheme. The simulations are performed in a square lattice of size $L\times L$, and the space is discretized as $L=N\Delta$. The grid spacing $\Delta$ is kept constant while varying the $N$ to change the system size. The initial condition for $\phi$ is sampled from randomly in [0.1,0.1][-0.1,0.1], and $\psi$ is set to zero.
@@
Figure 8: (a) EPR is plotted as a function of activity for various spatial discretizations $\Delta$ by fixing $L$. (b) Variation of EPR with spatial discretizations. The inset plot shows EPR per lattice cell. (c) Plot is same as (a) but by changing $L$ with fixed $\Delta$.
@@ Appendix D Derivation of the informatic entropy production rate
$\dot{\cal S}_{\psi}={+}\frac{1}{T}\int\langle L\psi\rangle\,\mathrm{d}{\bf\it r},\quad\dot{\cal S}_{\phi}=-\frac{1}{T}\int\left\langle\left(\dot{\phi}+{\bf\it v}\cdot\nabla\phi\right)\frac{\delta\cal F}{\delta\phi}\right\rangle\,\mathrm{d}{\bf\it r},$ (66)
@@
$T\dot{\cal S}_{\phi}$ $={-}\epsilon_{\alpha\beta}\int\langle(\partial_{\beta}\psi)(\partial_{\alpha}\phi)(dg/d\phi-\kappa\nabla^{2}\phi)\rangle\,\mathrm{d}{\bf\it r}$ (68)
$={-}\epsilon_{\alpha\beta}\int\langle(\partial_{\beta}\psi)(\partial_{\alpha}g-\kappa(\partial_{\alpha}\phi)\nabla^{2}\phi)\rangle\,\mathrm{d}{\bf\it r}$
$={+}\epsilon_{\alpha\beta}\int\langle\psi(\partial_{\beta}\partial_{\alpha}g-\kappa\partial_{\beta}((\partial_{\alpha}\phi)\nabla^{2}\phi))\rangle\,\mathrm{d}{\bf\it r},$
where we have integrated by parts. Given that $\epsilon_{\alpha\beta}\partial_{\alpha}\partial_{\beta}h(\phi)=0$ for an arbitrary function $h$ (since $\epsilon_{\alpha\beta}=-\epsilon_{\alpha\beta}$), we deduce
$T\dot{\cal S}_{\phi}={-}\kappa\epsilon_{\alpha\beta}\int\langle\psi(\partial_{\alpha}\phi)\nabla^{2}(\partial_{\beta}\phi)\rangle\,\mathrm{d}{\bf\it r}.$ (69)
@@
$T\dot{\cal S}_{\psi}={+}(\zeta+\kappa)\epsilon_{\alpha\beta}\int\langle\psi(\partial_{\alpha}\phi)\nabla^{2}(\partial_{\beta}\phi)\rangle\,\mathrm{d}{\bf\it r}.$ (70)
@@ Appendix F Independence of IEPR from boundary conditions
Appendix G Spatial distribution of the active force
Figure 10: Prediction of the qualitative localization of the IEPR derived from various force fields. The plots are in arbitrary units. (a,b) Exact force field and corresponding IEPR form numerical simulations. Panel (c): Analytic force field around an isolated $\frac{1}{2}$ defect (analog to active nematics, [74]). Note that the force is maximal at the center of the defect, contrary to (a). Consequently, while the mirror symmetry is qualitatively preserved, the IEPR [Panel (d)] is maximal at the center, and there is no separation between the two lobes of IEPR. Panel (e): Analytic ansatz for the force field around a $\frac{1}{2}$ defect, which qualitatively reproduces the symmetries of [Panel (a), inset] as well as the localization along the interface. The corresponding IEPR [Panel (f)] has two localized lobes consistent with the full solution shown in Panel (b).
In this Appendix, we comment on the relation between the local IEPR $\dot{\sigma}$ [Eq. (19)] and the distribution of the total force $f_{\alpha}\propto\partial_{\beta}Q_{\alpha\beta}$ [Eq. (8)] around individual defects by comparing numerical simulations with some analytical profiles. In the turbulent state, we consider a specific region with a $\frac{1}{2}$ defect, whose orientation is indicated with a green arrow [Fig. 10(a)]. The corresponding IEPR is maximal on the two sides of the defect [Fig. 10(b)], is approximately symmetric with respect to reflections across the defect’s orientation, and is comparatively lower near the origin. In what follows, we provide some arguments to reproduce the profiles of $\dot{\sigma}$ and $f_{\alpha}$ around the defect.
In Sec. 3.3.1, our analytic profile of the local IEPR $\dot{\sigma}$ around a point-like defect [Eq. (33a)] assumes that the defect is isolated and the core size is zero, namely the amplitude of the $Q$-tensor is constant everywhere except at the origin [C]. As a result, the corresponding force profile $|{\bf f}|$ is strongly localized at the defect core [Fig. 10(c)], and $\dot{\sigma}$ is maximal at the defect core with a mirror symmetry with respect to the defect orientation [Fig. 10(d)], in contrast with the numerical profiles [Figs. 10(a-b)].
We now show that the lobed-structure of Figs. 10(a-b) is related to the non-constant amplitude of the $Q$-tensor close to the defect core. We propose an ansatz for the force field $\hat{\mathbf{f}}(\mathbf{r})$ that qualitatively mimics the geometry of the numerical profile:
[…]
where we assume that the defect is at the origin of the coordinate system, $r=\sqrt{x^{2}+y^{2}}$, and $\theta=\arctan(y/x)$. The ansatz in Eq. (92) is localized along the interface, vanishes at the defect core, and has a mirror symmetry [Fig. 10(e)], as seen in the numerical solution [Fig. 10(a)]. The corresponding IEPR profile follow by using Eqs. (19) and (35): it features two well-separated lobes located symmetrically with respect to the defect orientation [Fig. 10(f)] in agreement with simulations [Fig. 10(b)].
In short, we have shown that, for an individual defect, the geometry of the IEPR profile $\dot{\sigma}$ is encoded in the details of the corresponding force profile $|{\bf f}|$ close to the defect core. For defect pairs [Sec. 3.3.2], the deviation between our analytics (that rely on point-defect assumption) and numerics is less severe [Figs. 5(d-g)]. In this case, the constructive interference of vorticities occurs along the line connecting the two defects centers: the profile $\dot{\sigma}$ is overwhelmingly dominated by these vorices, so the details close to defect core are largely irrevelant.